If bob is 1/3 of Stan age and their total age combine together is 100, what is each of their ages?
Answer by thesvw(77)

Let Stan age be S and Bobs age be B.
B = (1/3)*S
B + S = 100
3B = S
B + 3B = 100
4B = 100
B = 25
Bobs age = 25 years
Stans age = 75 years